Solve LOGa(7x+1)=LOGa(4x+16)
Answer by ankor@dixie-net.com(22740)

:
Solve: LOGa(7x+1)=LOGa(4x+16)
:
I see this is a couple days old so no doubt you have figured out that:
"When the logs are equal the numbers are equal"
Therefore:
7x+1 = 4x + 16
Solve:
7x - 4x = 16 -1
x = 15/3
x = 5
Substitute: 7(5) + 1 = 4(5) + 16
:
Obviously log(36) = log(36)